### v1

# Electrically switchable non-relativistic Zeeman spin splittings in collinear antiferromagnets

Abstract
Magnetic or electrical manipulation of electronic spin is elementary for spin-based logic, computing, and memory, where the latter is a low-power manipulation scheme. Rashba-like spin splittings stemming from spin-orbit interaction (SOI) enable electric-field manipulation of spin, but the relativistic SOI causes spin relaxations and yields dissipative transport of spin-encoded information. Recent works suggest the occurrence of electric-field switchable non-relativistic Zeeman spin splittings (NRZSSs) in collinear antiferromagnets — allowing for electrical manipulation of spin in the non-relativistic regime; yet, a theory elucidating the mechanisms for these NRZSSs and guiding the materials discovery remains missing. Here, we develop such a theory by analyzing the symmetries of spin point groups. We highlight the linear magnetoelectric and bilinear piezomagnetoelectric mechanisms for NRZSSs that depend linearly on electric field and are electrically switchable. First-principles calculations further confirm that LiMnPO4 and NaMnP showcase such NRZSSs. Our theory provides guidelines for discovering light-element collinear antiferromagnets with electrically switchable NRZSSs, which are promising for the design of high-performance spin-based devices.
Introduction — Spintronic devices (e.g., logic, computing, and memory) are built upon the manipulation of electronic spin in solids [1, 2, 3, 4, 5]. Such a manipulation can be achieved by magnetic field or electric field, where the latter offers fascinating low-power devices by reducing the “Joule heating” [6, 7, 8]. In principle, electric-field manipulation of spin is enabled by electrically switchable spin splittings [9, 10]. Because of the relativistic spin-orbit interaction (SOI), dielectrics and ferroelectrics naturally host electrically switchable Rashba-like spin splittings [11, 12, 13, 14, 15]. As a by-product, SOI causes the so-called D’yakonov-Perel’ spin relaxation as well [16, 17, 18], which yields the spin decoherence of electrons and prevents the long-range transport of the spin-encoded information [19, 18, 20, 21]. To suppress the D’yakonov-Perel’ spin relaxation, the proposal of electrically switchable spin splittings in the non-relativistic regime is of high necessity. Recent works suggest the possibility of achieving non-relativistic Zeeman spin splittings (NRZSSs) in collinear antiferromagnets that are switchable by an electric field [22, 23, 24, 25, 26, 27, 28, 29, 30]. Yet, mechanisms for such NRZSSs are elusive and selection rules for the corresponding materials discovery remain to be explored.
To address the aforementioned issues, we analyze the symmetries of spin point groups (SPGs) that are associated with collinear magnets in the non-relativistic limit. This allows us to develop a theory on electrically switchable NRZSSs in collinear antiferromagnets, and reveal the magnetoelectric and piezomagnetoelectric mechanisms for this phenomenon. Our theory provides theoretical guidelines for screening collinear antiferromagnets with electrically switchable NRZSSs. This yields the first-principles predictions that LiMnPO4 and NaMnP are two candidates with linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs, respectively.
[…]
For all the $[R_{s}\left|\right|R_{l}]$ symmetry operations in ferromagnetic or ferrimagnetic SPGs, $R_{s}s_{\chi}=s_{\chi}$ is always true and $\varepsilon(s_{\chi})=\varepsilon(-s_{\chi})$ is not necessarily enforced [36, 37]. This means that non-relativistic collinear ferromagnets or ferrimagnets enable the spontaneous NRZSSs (see e.g., Refs. [33, 34, 35]). Unlike those in ferromagnetic or ferrimagnetic SPGs, the $[R_{s}\left|\right|R_{l}]$ symmetry operations in an antiferromagnetic SPG can be partitioned into two subsets [36, 37, 34, 35], namely,
$L_{+}=\{[R_{s}^{+}\left|\right|R_{l}^{+}]:R_{s}^{+}s_{\chi}=s_{\chi}\},$ (1)
and
$L_{-}=\{[R_{s}^{-}\left|\right|R_{l}^{-}]:R_{s}^{-}s_{\chi}=-s_{\chi}\}.$ (2)
The $L_{+}$ subset is a subgroup of the antiferromagnetic SPG, and the symmetry operations therein have no restriction on $\varepsilon(s_{\chi})$. In contrast, the symmetry operations in $L_{-}$ suggest the $\varepsilon(s_{\chi})=\varepsilon(-s_{\chi})$ degeneracy, and prevent the spontaneous NRZSSs [33, 34, 35]. On the other hand, the application of external stimuli may induce NRZSSs in non-relativistic collinear antiferromagnets. This is achieved by that such stimuli break all the symmetry operations in the $L_{-}$ set.
Electric-field induced NRZSSs — We concentrate on the NRZSSs in nonpolar collinear antiferromagnets that are driven by external electric field. To illustrate our basic logic, we work with an antiferromagnet whose SPG is the union of $L_{+}$ and $L_{-}$ sets [see Eqs. (1) and (2)]. The application of electric field $E_{\alpha}$ along $\alpha$ direction breaks the symmetries of the antiferromagnet ($\alpha$ being $x$, $y$, or $z$). This reduces the SPG of the antiferromagnet to the union of $L_{+}^{\alpha}$ and $L_{-}^{\alpha}$ sets, where
$\begin{split}L_{+}^{\alpha}=\{[R_{s}^{+}\left|\right|R_{l}^{+}]\in L_{+}:R_{l}^{+}E_{\alpha}=E_{\alpha}\},\\ L_{-}^{\alpha}=\{[R_{s}^{-}\left|\right|R_{l}^{-}]\in L_{-}:R_{l}^{-}E_{\alpha}=E_{\alpha}\},\end{split}$ (3)
and the “$\alpha$” superscript labels the symmetry reduction caused by the $E_{\alpha}$ electric field  [^3] When $E_{\alpha}$ is not compatible with the spatial $R_{l}$ operation (i.e., $R_{l}E_{\alpha}\neq E_{\alpha}$), the application of $E_{\alpha}$ break the $[R_{s}\left|\right|R_{l}]$ symmetry operation.. When $L_{-}^{\alpha}$ becomes empty, there will be no symmetry operations that enforce the $\varepsilon(s_{\chi})=\varepsilon(-s_{\chi})$ degeneracy — enabling the electric-field induced NRZSSs. The $E_{\alpha}$-field-induced NRZSSs are described by the effective Hamiltonian
$\begin{split}\mathcal{H}=\lambda_{1}E_{\alpha}\sigma_{\chi}+\lambda_{2}E_{\alpha}^{2}\sigma_{\chi}+\cdots,\end{split}$ (4)
with $\lambda_{n}$ ($n=1,2,\cdots$) characterizing the $n_{\mathrm{th}}$-order response of the Zeeman splitting to the $E_{\alpha}$ field, and $\sigma_{\chi}$ representing the Pauli matrix along the $\chi$ direction. Eq. (4) is reminiscent of the magnetoelectric coupling, and this motivates us to name the electric-field induced NRZSSs as the magnetoelectric NRZSSs.
The $L_{-}^{\alpha}$ set associated with an SPG may be not empty, and a material with this SPG does not host the $E_{\alpha}$-induced NRZSSs. In such a case, we can grow this material on an appropriate substrate so that an epitaxial $\eta_{\beta\gamma}$ strain is exerted to this material. The $\eta_{\beta\gamma}$ strain might modify the symmetry of the material, and reduce the $L_{-}$ set to $L_{-}^{\beta\gamma}=\{[R_{s}^{-}\left|\right|R_{l}^{-}]\in L_{-}:R_{l}^{-}\eta_{\beta\gamma}=\eta_{\beta\gamma}\}$. Applying the $E_{\alpha}$ electric field to the strained material further reduces $L_{-}^{\beta\gamma}$ to
$L_{-}^{\alpha\beta\gamma}=\{[R_{s}^{-}\left|\right|R_{l}^{-}]\in L_{-}^{\beta\gamma}:R_{l}^{-}E_{\alpha}=E_{\alpha}\}.$ (5)
Provided that $L_{-}^{\alpha\beta\gamma}$ is empty but $L_{-}^{\beta\gamma}$ is not, the NRZSSs are created via the cooperation of $E_{\alpha}$ and $\eta_{\beta\gamma}$  [^4] When $L_{-}^{\beta\gamma}$ is empty, the NRZSSs can be driven by the $\eta_{\beta\gamma}$ strain solely, without involving the $E_{\alpha}$ electric field.. Phenomenologically, such NRZSSs are described by the effective Hamiltonian
$\mathcal{H}^{\prime}=\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}+\lambda^{\prime}_{2}E_{\alpha}^{2}\eta_{\beta\gamma}\sigma_{\chi}+\lambda^{\prime}_{3}E_{\alpha}\eta_{\beta\gamma}^{2}\sigma_{\chi}+\cdots,$ (6)
with $\lambda^{\prime}_{n}$ ($n=1,2,\cdots$) being the coupling coefficients. Eq. (6) motivates us to name this type of NRZSSs as piezomagnetoelectric NRZSSs.
Linear magnetoelectric and bilinear piezomagnetoelectric NRZSSs — We continue to carry out symmetry analysis and extract the SPGs that host the electric-field switchable NRZSSs, following the logics mentioned in the previous section. Our detailed analysis can be found in Sections I and II of the Supplementary Material (SM), where the SPGs hosting magnetoelectric and piezomagnetoelectric NRZSSs are summarized in Tables S6 and S9 therein. As shown in Eqs. (4) and (6), the $E_{\alpha}$-driven magnetoelectric or piezomagnetoelectric NRZSSs may be rooted in $E_{\alpha}\sigma_{\chi}$, $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$, $E_{\alpha}^{2}\sigma_{\chi}$, $E_{\alpha}\eta_{\beta\gamma}^{2}\sigma_{\chi}$, or other higher-order couplings (see Tables S11 and S12 of the SM). Here, we focus on the predominant electric-field induced NRZSSs which include linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs. In Tables 1 and 2, we collect the nonpolar collinear SPGs that enable these two types of NRZSSs.
The linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs are characterized by $\lambda_{1}E_{\alpha}\sigma_{\chi}$ and $\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$, respectively. At the Brillouin zone center, the electronic eigen energies are $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\pm\lambda_{1}E_{\alpha}$ or $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\pm\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}$, with $\varepsilon_{0}$ being an energy independent of spin. When changing the electric field from $E_{\alpha}$ to $-E_{\alpha}$, the eigen energies become $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\mp\lambda_{1}E_{\alpha}$ or $\varepsilon(\pm s_{\chi})=\varepsilon_{0}\mp\lambda^{\prime}_{1}E_{\alpha}\eta_{\beta\gamma}$. This means that both linear magnetoelectric NRZSSs and bilinear piezomagnetoelectric NRZSSs are electric-field switchable.
Table 1: The nonpolar collinear SPGs that enable linear magnetoelectric NRZSSs. For an SPG associated with the $E_{\alpha}\sigma_{\chi}$ coupling, a $\circ$” symbol is filled in the corresponding entry; Such an entry is endowed with a “--” symbol if the $E_{\alpha}\sigma_{\chi}$ coupling is not symmetrically allowed. For short, we relabel each collinear $G^{\infty m}1$ SPG as $G$ by omitting its ${}^{\infty m}1$ part. As an example, the ${}^{1}m^{1}m^{\bar{1}}m^{\infty m}1$ SPG (abbreviated as ${}^{1}m^{1}m^{\bar{1}}m$) hosts the $E_{z}$-driven linear magnetoelectric NRZSSs via the $E_{z}\sigma_{\chi}$ coupling, but does not accommodate the $E_{x}$- or $E_{y}$-driven linear magnetoelectric NRZSSs.
SPGs $E_{x}$ $E_{y}$ $E_{z}$ SPGs $E_{x}$ $E_{y}$ $E_{z}$
${}^{\bar{1}}\bar{1}$circ $\circ$ $\circ$ ${}^{\bar{1}}2/^{1}m$circ -- $\circ$
${}^{1}2/^{\bar{1}}m$- $\circ$ -- ${}^{\bar{1}}2^{\bar{1}}2^{1}2$- -- $\circ$
${}^{\bar{1}}2^{1}2^{\bar{1}}2$- $\circ$ -- ${}^{1}2^{\bar{1}}2^{\bar{1}}2$circ -- --
${}^{\bar{1}}m^{1}m^{1}m$circ -- -- ${}^{1}m^{\bar{1}}m^{1}m$- $\circ$ --
${}^{1}m^{1}m^{\bar{1}}m$- -- $\circ$ ${}^{\bar{1}}\bar{4}$- -- $\circ$
${}^{1}4/^{\bar{1}}m$- -- $\circ$ ${}^{1}4^{\bar{1}}2^{\bar{1}}2$- -- $\circ$
${}^{\bar{1}}\bar{4}^{\bar{1}}2^{1}m$- -- $\circ$ ${}^{1}4/^{\bar{1}}m^{1}m^{1}m$- -- $\circ$
${}^{\bar{1}}\bar{3}$- -- $\circ$ ${}^{1}3^{\bar{1}}2$- -- $\circ$
${}^{\bar{1}}\bar{3}^{1}m$- -- $\circ$ ${}^{\bar{1}}\bar{6}$- -- $\circ$
${}^{1}6/^{\bar{1}}m$- -- $\circ$ ${}^{1}6^{\bar{1}}2^{\bar{1}}2$- -- $\circ$
${}^{\bar{1}}\bar{6}^{1}m^{\bar{1}}2$- -- $\circ$ ${}^{1}6/^{\bar{1}}m^{1}m^{1}m$- -- $\circ$
Table 2: The nonpolar collinear SPGs that enable the bilinear piezomagnetoelectric NRZSSs. The $\eta_{xx}$, $\eta_{yy}$, $\eta_{zz}$, $\eta_{xy}$, $\eta_{yz}$, and $\eta_{zx}$ strains are represented by “1”, “2”, “3”, “4”, “5”, and “6” numbers, respectively. For an SPG associated with the $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$ coupling, the number representing the $\eta_{\beta\gamma}$ strain is filled in the corresponding entry. If an SPG enables $E_{\alpha}\eta_{\beta\gamma}\sigma_{\chi}$ and $E_{\alpha}\eta_{\beta^{\prime}\gamma^{\prime}}\sigma_{\chi}$ couplings, two numbers representing the $\eta_{\beta\gamma}$ and $\eta_{\beta^{\prime}\gamma^{\prime}}$ strains are filled in the corresponding entry, and so forth. The entry with “--” indicates that the corresponding bilinear piezomagnetoelectric NRZSSs are symmetrically forbidden. In this table, we relabel each collinear $G^{\infty m}1$ SPG as $G$ by omitting its ${}^{\infty m}1$ part. As an example, the ${}^{\bar{1}}\bar{3}^{1}m^{\infty m}1$ SPG (abbreviated as ${}^{\bar{1}}\bar{3}^{1}m$ in this table) hosts the $E_{x}$-driven bilinear piezomagnetoelectric NRZSSs via $E_{x}\eta_{xy}\sigma_{\chi}$ or $E_{x}\eta_{zx}\sigma_{\chi}$ couplings, but does not accommodate the $E_{y}$- and $E_{z}$-driven bilinear piezomagnetoelectric NRZSSs.
SPGs $E_{x}$ $E_{y}$ $E_{z}$ SPGs $E_{x}$ $E_{y}$ $E_{z}$
${}^{\bar{1}}2/^{1}m$- $45$- ${}^{1}2/^{\bar{1}}m$ $45$- $45$
${}^{\bar{1}}2^{\bar{1}}2^{1}2$ $6$ $5$- ${}^{\bar{1}}2^{1}2^{\bar{1}}2$ $4$- $5$
${}^{1}2^{\bar{1}}2^{\bar{1}}2$- $4$ $6$ ${}^{\bar{1}}m^{1}m^{1}m$- $4$ $6$
${}^{1}m^{\bar{1}}m^{1}m$ $4$- $5$ ${}^{1}m^{1}m^{\bar{1}}m$ $6$ $5$-
${}^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ $5$ $6$ $4$ ${}^{1}4/^{\bar{1}}m$ $56$ $56$-
${}^{\bar{1}}4/^{\bar{1}}m$ $56$ $56$ $124$ ${}^{\bar{1}}4^{1}2^{\bar{1}}2$- -- $4$
${}^{\bar{1}}4^{\bar{1}}2^{1}2$ $6$ $5$ $12$ ${}^{1}4^{\bar{1}}2^{\bar{1}}2$ $6$ $5$-
${}^{\bar{1}}4^{\bar{1}}2^{1}m$ $6$ $5$- ${}^{1}4^{\bar{1}}2^{\bar{1}}m$ $6$ $5$ $12$
${}^{1}4/^{\bar{1}}m^{1}m^{1}m$ $6$ $5$- ${}^{\bar{1}}4/^{\bar{1}}m^{\bar{1}}m^{1}m$ $5$ $6$ $4$
${}^{\bar{1}}4/^{\bar{1}}m^{1}m^{\bar{1}}m$ $6$ $5$ $12$ ${}^{1}4/^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ $5$ $6$-
${}^{\bar{1}}\bar{3}^{1}m$ $46$- -- ${}^{\bar{1}}\bar{3}^{\bar{1}}m$- $46$-
${}^{1}6/^{\bar{1}}m$ $56$ $56$- ${}^{\bar{1}}6^{\bar{1}}2^{1}2$ $4$- --
${}^{\bar{1}}6^{1}2^{\bar{1}}2$- $4$- ${}^{1}6^{\bar{1}}2^{\bar{1}}2$ $6$ $5$-
${}^{\bar{1}}\bar{6}^{\bar{1}}m^{1}2$ $5$- -- ${}^{\bar{1}}\bar{6}^{1}m^{\bar{1}}2$ $6$- --
${}^{1}\bar{6}/^{\bar{1}}m^{1}m^{1}m$ $6$ $5$- ${}^{\bar{1}}\bar{6}/^{1}m^{\bar{1}}m^{1}m$- $4$-
${}^{\bar{1}}\bar{6}/^{1}m^{1}m^{\bar{1}}m$ $4$- -- ${}^{1}\bar{6}/^{\bar{1}}m^{\bar{1}}m^{\bar{1}}m$ $5$ $6$-
${}^{\bar{1}}m^{\bar{1}}\bar{3}$ $5$ $6$ $4$ ${}^{\bar{1}}4^{1}3^{\bar{1}}2$ $5$ $6$ $4$
${}^{\bar{1}}m^{\bar{1}}\bar{3}^{1}m$ $5$ $6$ $4$
Materials with electric-field switchable NRZSSs — Assisted by the MAGNDATA database [43, 44], and guided by Tables 1 and 2, we explore materials with linear magnetoelectric NRZSSs or bilinear piezomagnetoelectric NRZSSs. We search for nonpolar collinear antiferromagnets composed of light elements  [^5] The strength of SOI basically depends on the atomic mass. Materials composed of heavy elements (e.g., Bi, Sb, and Pt) likely showcase sizable SOI [50, 51, 52]., identify the SPGs for such antiferromagnets, and discover LiMnPO4 and NaMnP as two representatives with electric-field switchable NRZSSs. The antiferromagnetic LiMnPO4 has a Néel temperature of 34 K [46, 47] and an SPG of ${}^{\bar{1}}m^{1}m^{1}m^{\infty m}1$. According to Table 1, the ${}^{\bar{1}}m^{1}m^{1}m^{\infty m}1$ SPG enables linear magnetoelectric NRZSSs via the $E_{x}\sigma_{\chi}$ coupling. The NaMnP material is a room-temperature antiferromagnet with a Néel temperature of 560 K [48, 49]. The SPG of NaMnP is ${}^{\bar{1}}4/^{\bar{1}}m^{\bar{1}}m^{1}m^{\infty m}1$, and this group hosts the piezomagnetoelectric NRZSSs induced by the combination of $\eta_{xy}$ and $E_{z}$ (see the $\eta_{xy}E_{z}\sigma_{\chi}$ coupling in Table 2). To validate our analysis, we compute the band structures for LiMnPO4 and NaMnP antiferromagnets. First of all, we neglect the SOI and carry out collinear magnetism calculations. This treats the magnetic moments of Mn in LiMnPO4 and NaMnP as scalar quantities (see Fig. 1). On this condition, the electronic spin degree of freedom appears as unmixed spin-up $+s$ and spin-down $-s$ states, where $\chi$ in $s_{\chi}$ is omitted at the collinear magnetism level.
Figure 1: Collinear magnetic structures for (a) LiMnPO4 and (b) NaMnP antiferromagnets, where only the Mn sublattices are displayed. The cyan and pink spheres denote Mn ions with magnetic moments along $+\bm{\chi}$ and $-\bm{\chi}$ directions, respectively. In panels (a) and (b), the boxes enclosed by the black dashed lines represent the cells of LiMnPO4 and NaMnP employed in our simulations. The primitive cell of NaMnP is sketched by the box with pink solid lines in panel (b).
As shown in Fig. S2 of the SM, both the conduction band minimum (CBM) of LiMnPO4 and the valence band maximum (VBM) of NaMnP locate at the Brillouin zone center. We shall focus on such two band edges in our following discussion. Under null electric field, the spin-up $\varepsilon_{\mathrm{C}}(+s)$ and spin-down $\varepsilon_{\mathrm{C}}(-s)$ energy levels of LiMnPO4 are degenerate at the CBM. Applying an electric field along the $x$ direction splits the $\varepsilon_{\mathrm{C}}(+s)$ and $\varepsilon_{\mathrm{C}}(-s)$ energy levels. To be accurate, the $\Delta_{1}=\varepsilon_{\mathrm{C}}(+s)-\varepsilon_{\mathrm{C}}(-s)$ splittings are $\mp 23$ meV for $E_{x}=\pm 6$ MV/cm [see Figs. 2(a) and 2(b)]. For unstrained NaMnP under null electric field, the spin-up $\varepsilon_{\mathrm{V}}(+s)$ and spin-down $\varepsilon_{\mathrm{V}}(-s)$ energy levels at the VBM are degenerated as well. In the presence of epitaxial strain $\eta_{xy}=4\%$, the $\Delta_{2}=\varepsilon_{\mathrm{V}}(+s)-\varepsilon_{\mathrm{V}}(-s)$ splittings of ±$\pm 26$ meV can be created by applying $E_{z}$ of ±$\pm 6$ MV/cm. Figure 3 shows the $\Delta_{1}$ splittings in LiMnPO4 as a function of $E_{x}$ and the $\Delta_{2}$ splittings in NaMnP as a function of $E_{z}$ ($\eta_{xy}=4\%$). This confirms that both the magnetoelectric NRZSSs in LiMnPO4 (CBM) and piezomagnetoelectric NRZSSs (VBM) in NaMnP depend linearly on the corresponding electric fields, and are electrically switchable.
To complete this section, we check the relativistic effect on the electric-field induced Zeeman-type spin splittings in LiMnPO4 and NaMnP. We calculate the band structures for LiMnPO4 and NaMnP at noncollinear magnetism level (with SOI). This treats the magnetic moments of Mn as vectors along ±$\pm\bm{\chi}$ directions ($\bm{\chi}=\mathbf{x},\mathbf{y},\mathbf{z}$). Because of the SOI, the spin-up and spin-down eigen states are not well defined in LiMnPO4 and NaMnP. Instead, the electronic spin degree of freedom is characterized by the spin angular momentum vector, whose predominant component is along $+\bm{\chi}$ or $-\bm{\chi}$ direction. In this case, we may use $\Delta^{\prime}_{1}=\varepsilon_{C}(+s_{\chi})-\varepsilon_{C}(-s_{\chi})$ and $\Delta^{\prime}_{2}=\varepsilon_{V}(+s_{\chi})-\varepsilon_{V}(-s_{\chi})$ to describe the Zeeman-type spin splittings associated with the CBM of LiMnPO4 and the VBM of NaMnP, respectively (see the previous paragraph for the definitions of $\Delta_{1}$ and $\Delta_{2}$). In Fig. S3 of the SM, we show the band structures with Zeeman-type spin splittings for LiMnPO4 and NaMnP, calculated at the relativistic level. Basically, the inclusion of relativistic corrections does not affect the magnitude of the Zeeman-type spin splittings in LiMnPO4 and NaMnP.
Figure 2: The electric-field switchable NRZSSs in LiMnPO4 and NaMnP. Panels (a) and (b) are magnetoelectric NRZSSs in LiMnPO4 driven by $E_{x}$, while panels (c) and (d) are piezomagnetoelectric NRZSSs in NaMnP driven by $E_{z}$ combined with $\eta_{xy}=4\%$. The red (blue) solid lines correspond to the spin-up (spin-down) states, and $\varepsilon_{F}$ denotes the fermi level.
Figure 3: The linear magnetoelectric $\Delta_{1}$ splittings in LiMnPO4 and bilinear piezomagnetoelectric $\Delta_{2}$ splittings in NaMnP. Purple circles show $\Delta_{1}$ in LiMnPO4 as a function of $E_{x}$; Blue squares show $\Delta_{2}$ in NaMnP as a function of $E_{z}$ (under a fixed strain of $\eta_{xy}=4\%$). Purple and blue lines display the corresponding linear regression results.
Summary and outlook — By analyzing the symmetries of spin point groups, we develop a theory that describes the NRZSSs in collinear antiferromagnets driven by electric field. In particular, we identify the linear magnetoelectric and bilinear piezomagnetoelectric mechanisms, both of which are associated with NRZSSs depending linearly on the applied electric field. The reversal of electric field flips the electronic spin states, yielding the electric-field switchable NRZSSs. Our central results are summarized in Tables 1 and 2, and this guides the first-principles predictions of linear magnetoelectric NRZSSs in LiMnPO4 and bilinear piezomagnetoelectric NRZSSs in NaMnP.
Our theory provides guidelines for the discovery of light-element collinear antiferromagnets with linear magnetoelectric or bilinear piezomagnetoelectric NRZSSs. The NRZSSs therein resemble Zeeman spin splittings in conventional ferromagnets or ferrimagnets, but are more attractive due to the merits of antiferromagnets (e.g., ultralow stray field and fast spin dynamics [1, 2]). Moreover, light-element materials exhibit weak SOIs that result in suppressed spin relaxations and enhanced spin life times, and the electrically switchable NRZSSs enable the manipulation of spin in an energy-efficient electrical manner. These features highlight the potential of such light-element collinear antiferromagnets in designing high-performance spin-based information devices.
Acknowledgements — The authors acknowledge the support from the National Natural Science Foundation of China (Grants No. 12274174, No. 52288102, No. 52090024, and No. 12034009). L.B. thanks the Vannevar Bush Faculty Fellowship (VBFF) grant No. N00014-20-1-2834 from the Department of Defense and award No. DMR-1906383 from the National Science Foundation AMASE-i Program (MonArk NSF Quantum Foundry). L.Y. thanks the support from high-performance computing center of Jilin University. H.J.Z. acknowledges the support from “Xiaomi YoungScholar” Project.